Lesco Chemical is considering two processes for making a cationic polymer. Process A has a first cost of $120,000 and an annual operating cost (AOC) of $55,000 per year.

The first cost of process B is $165,000. If both processes will be adequate for 7 years and the rate of return on the increment between the alternatives is 25%, what is the amount of the AOC for process B?

The amount of the AOC for process B is $__________ .
Answer #1

Incremental initial cost over process A=-(165000-120000)=-45000

Incremental AOC over process A=-(AOC-55000)=-AOC+55000

Incremental Rate of return=i*=25%

Let us find incremental cash flows

Incremental NPW=-45000+(-AOC+55000)*(P/A,25%,7)

It should be equal to zero at 25% incremental ROR

-45000+(-AOC+55000)*(P/A,25%,7)=0

-45000-AOC*(P/A,25%,7)+55000*(P/A,25%,7)=0

We can get the value of interest factor from tables with 4-decimal places or 6-decimal places.

I shall consider the value with 6-decimal places for better accuracy.

(P/A,0.25,7)=3.161139

I am showing formula to get interest factor. Same can also be obtained from tables

Let us calculate AOC.

AOC is cost, it will be negative.

AOC=-40765